9. Zach took $200,000 out of the bank and used it to start his new cookie business. The bank account pays 3 percent interest per year. During the first year of his business, Zach sold 17,000 boxes of cookies for $5 per box. Also, during the first year, the cookie business incurred costs amounting to $50,000.

a. What was Zach’s accounting profit for the year?

b. What was Zach’s economic profit for the year?

c. Should Zach continue with his cookie business?

d. A key difference between accountants and economists is their different treatment of the cost of capital. Does this difference result in an accountant’s estimate of total costs being higher or lower than an economist’s estimate? Explain your answer.
Answer #1

a)

Accounting profit = Total Revenue - Explicit Cost.

Total Revenue = 17,000*5

= 85,000

Explicit Cost = 50 000

Profit = 85000 - 50 000

= 35,000

b)

Economic Profit = Total Revenue (TR) - Total cost including implicit cost.

Total Revenue = 85, 000

Explicit cost = 50,000

Implicit cost = 200000*0.03

= 6000

Economic Profit = 35000 - 6000

= 29000

c)

Both accounting and economist profits are positive, so Zach would continue to produce.

d)

Accountant takes into account only explicit cost while he does not consider opportunity cost or implicit cost. Hence, While economists consider both types of costs, explicit and implicit. Hence,. cost estimated by the economists is usually greater relative to accountant